OpenStudy (anonymous):

Suppose that g is an integrable periodic function with period T. Show that for any values of a and b that the definite integral from a to a+T of g(x) is equal to the definite integral from b to b+T of g(x)

OpenStudy (anonymous):

\[\int\limits_{a}^{a+T}g(x)dx = \int\limits_{b}^{b+T}g(x)dx\]

OpenStudy (anonymous):

alright, here we go~ WLOG, assume a < b \[\int\limits_{a}^{a+T}g(x)dx = \int\limits_{a}^{b+T}g(x)dx - \int\limits_{a+T}^{b+T}g(x)dx = \int\limits_{a}^{b}g(x)dx+\int\limits_{b}^{b+T}g(x)dx-\int\limits_{a+T}^{b+T}g(x)dx\] but, because the function is periodic: \[\int\limits_{a+T}^{b+T}g(x)dx = \int\limits_{a}^{b}g(x)dx\] So now we have: \[\int\limits_{a}^{b}g(x)dx+\int\limits_{b}^{b+T}g(x)dx-\int\limits_{a}^{b}g(x)dx = \int\limits_{b}^{b+T}g(x)dx\]

OpenStudy (anonymous):

whoops, that should be: \[\int\limits_{a+T}^{b+T}g(x)dx\] in the second equation line.

OpenStudy (anonymous):

what!? why wont it put my a in the bottom limit lol <.<

OpenStudy (anonymous):

you got this. i knew it!

OpenStudy (anonymous):

one more time: \[\int\limits_{a+T}^{b+T}g(x)dx\]

OpenStudy (anonymous):

<.< its not me!

OpenStudy (anonymous):

\[\int_{a+T}^{b+T}g(x)dx\]

OpenStudy (anonymous):

the hell >.>

OpenStudy (anonymous):

is that what you want? i have only done this using \[g(x+t)\]

OpenStudy (anonymous):

lolol, well, thats what should be in my second equation line, what satellite posted.

OpenStudy (anonymous):

yes thats what i wanted lol

OpenStudy (anonymous):

Thanks, so I am guessing I just need to show this for a = b and a > b using similar kinds of arguments?

OpenStudy (anonymous):

Whoops scratch a = b....

OpenStudy (anonymous):

really scratch that!

OpenStudy (anonymous):

and no you don't have to do it again. that is what wlog means

OpenStudy (anonymous):

WLOG?

OpenStudy (anonymous):

without loss of generality. one of them has to be smaller, so it might as well be a.

OpenStudy (anonymous):

oh gotcha thx